Find: (v) 0.05 × 7 (vi) 211.02 × 4 (vii) 2 × 0.86 nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Find:
 

(v) 0.05 × 7

(vi) 211.02 × 4

(vii) 2 × 0.86

nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(v) 0.05 × 7

Solution:- We have,

= (5/100) × 7

= (35/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 0.35

(vi) 211.02 × 4

Solution:-
We have,

= (21102/100) × 4

= (84408/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 844.08

(vii) 2 × 0.86

Solution:- We have,

= (2) × (86/100)

= (172/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 1.72​

Thanks